In a class of 35 students, 19 take history and 12 take economics. if 5 take both subjects, how many take neither?
Answer by laoman(51)

This is a simple set question dear,
Given
Total no. of studenta(Universal Set) = 35
No. of students that take history = 19
No. of students that take economics = 12
No. of students that take both = 5
Now we need to determine,
No. of students that take history DISABLED_event_only= 19 - 5 = 14
No. of students that take economics DISABLED_event_only= 12 - 5 = 7
Ok, after knowing these,
we also know that
Total no. of stds = no. of stds taking history only + no. of stds taking econs only + no. of std taking both + no. of stds taking neither(which we are looking for)
Therefore, 35 = 14+7+5+x
where is the no. of stds taking neither
Solving the equation
35-14-7-5 = x
x= 9
Therefore, no. of stds taking neither subject is 9
Note that they are still students, just that they dont get involved in the two subjects